OpenStudy (idealist10):

Find a second solution y2 for (x-1)y"-xy'+y=0; y1=e^x that isn't a constant multiple of the solution y1.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Using reduction of order, we want to find a second solution \(y_2=uy_1\), where \(u,y_1,y_2\) are all functions of \(x\). \[y_2=uy_1\\ {y_2}'=u'y_1+u{y_1}'\\ {y_2}''=u''y_1+2u'{y_1}'+u{y_1}''\] With \(y_1=e^x\), you have \[y_2=ue^x\\ {y_2}'=(u'+u)e^x\\ {y_2}''=(u''+2u'+u)e^x\] Subbing into the ODE, you have \[\begin{align*} 0&=(x-1)y''-xy'+y\\\\ 0&=\bigg((x-1)(u''+2u'+u)-x(u'+u)+u\bigg)e^x\\\\ 0&=(x-1)u''+(2x-2)u'+(x-1)u-x(u'+u)+u\\\\ 0&=(x-1)u''+(x-2)u'\end{align*}\] Substituting \(t=u'\) gives you an equation linear in \(t\): \[0=(x-1)t'+(x-2)t\]

OpenStudy (idealist10):

So (x-1)t'=-(x-2)t?

OpenStudy (anonymous):

Right, you can solve this by separation of variables, or find an integrating factor. Your choice.

OpenStudy (idealist10):

But I'm not getting the answer in the book. \[t'=-\frac{ x-2 }{ x-1 }t\]

OpenStudy (idealist10):

I got \[t=C(x-1)e^ {-x}\]

OpenStudy (idealist10):

And \[u'=Ce^ {-x}(x-1)\]

OpenStudy (idealist10):

\[\frac{ u }{ c }=-xe^ {-x}+C\]

OpenStudy (idealist10):

The answer in the book is y2=x.

OpenStudy (anonymous):

\[t'=-\frac{ x-2 }{ x-1 }t~~\implies~~\dfrac{dt}{t}=-\frac{x-2}{x-1}\,dx\] Integrate both sides and back-substitute, \[\begin{align*} \int\dfrac{dt}{t}&=-\int\frac{x-2}{x-1}\,dx\\\\ \ln t&=-\int\frac{x-1-1}{x-1}\,dx\\\\ &=\int\left(\frac{1}{x-1}-1\right)\,dx\\\\ &=\ln(x-1)-x+C_1\\\\ t&=C_1(x-1)e^{-x}\\\\ u'&=C_1(x-1)e^{-x}\\\\ u&=-C_1xe^{-x}+C_2\\\\ ue^x&=-C_1x+C_2e^x\\\\ \implies y_2&=x\end{align*}\]

OpenStudy (idealist10):

I see my mistake now. :) Thanks a lot!